I'm learning absolute value functions and graphs and my book threw these harder equations in without show me how to do these more complex ones. I don't have and trouble with ones in the form y=|mx+b|+c. Please explain the steps to me that is more important than answers.
y=|3x|-x|1/3|
y=x-|2x|
y=1/2|x|+4|x-1|
y=|x+1|+|x|

Answer by Alan3354(69443)

I'm learning absolute value functions and graphs and my book threw these harder equations in without show me how to do these more complex ones. I don't have and trouble with ones in the form y=|mx+b|+c. Please explain the steps to me that is more important than answers.
------------
The only step is to pick values for x, find y, plot the points and draw a line thru the points.
It's all straight lines.
============================
y=|3x|-x|1/3|
|1/3| = 1/3
=========================
y=x-|2x|
y=1/2|x|+4|x-1|
y=|x+1|+|x|